How much heat is required to change 10g of ice at exactly 0°C to steam at 100°C ?


1
Expert's answer
2022-04-04T17:09:42-0400
"Q=Q_1+Q_2+Q_3\\
=\lambda m+cm\Delta T+Lm"

"Q=334*10+4.184*10*100+2257*10\\
=30094\:\rm J"